OpenStudy (anonymous):

I'm pretty bad at math, and I need some help with this question Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain The graph is a straight line, starting at Y-300, and going southwest in a straight line to X 450

OpenStudy (anonymous):

This is the graph

OpenStudy (needhelpwithk12):

well I am with you on the math thing but i might be able to figure it out.

OpenStudy (anonymous):

That would be amazing!

OpenStudy (needhelpwithk12):

ok I will try

OpenStudy (sleepyhead314):

Hello @Venom1954 and Welcome to OpenStudy! :) we can try using the equation of \(y = mx + b\) where \(m = slope\) and \(b = y-intercept\) In this case, they have given us that the \(y-intercept = 300\) so we just need to find the slope slope = rise divided by run = (high minus low) / (left minus right) = (300 - 0) / (0 - 450) = negative 300/450 = [simplify] once you get that, we can plug in m = slope and y intercept = 300 into the equation y = mx + b This should get you to your answer ^_^ If you are stuck or still confused, feel free to tag me with @sleepyhead314 :)